OpenStudy (anonymous):

x^(3/5)^(2/3) simplified

OpenStudy (anonymous):

i attempted to use the equation button but it had gotten mixed up.

OpenStudy (abb0t):

\[(x^{\frac{ 3 }{ 5} })^{\frac{ 2 }{ 3 }}\] correct?

OpenStudy (anonymous):

yes indeed.

OpenStudy (abb0t):

Just multiply the exponents and simplify. Remember this: \[(a^b)^c = a^{bc} \] However, \[(a^b )(a^c) = a^{b + c}\]

OpenStudy (abb0t):

Also: \[(a^{\frac{ c }{ L }})^{\frac{ d }{ Y }} = a^{\frac{ cd }{ LY }}\]

OpenStudy (anonymous):

please just the answer.